L13) A circuit has an emf of 12 V and internal resistance of 0.1 ohm. The seat of emf is placed across a resistor array that has two sets of three 5 ohm resistors in parallel. These two sets are in series with each other, and with an additional 2 ohm resistor. a) What is the current thru the 2 ohm resistor? b) After 2 years of use, the internal resistance grows to 0.5 ohms. What is the current thru the 2 ohm resistor in this case?
Answer #1

(a) Equivalent resistance of 03 nos. of 5 ohm resistor = 5/3 = 1.67 ohm

these two sets are in series.

So, equivalent resistance = 2*1.67 = 3.34 ohm

So, the total resistance across the circuit, R = 0.1 + 3.34 + 2 = 5.44 OHM

So the current through 2 ohm resistor = 12/5.44 = 2.21 A

(b) Total resistance, R = 0.5 + 3.34 + 2 = 5.84 OHM

So, the current, I = 12/5.84 = 2.05 A